# Non-minimal light-curvature couplings and black-hole imaging

### v1

@@ IV.1 Background on lensing bands
To explore how the non-minimal coupling $\alpha$ affects VLBI observations of supermassive black holes, we leverage the concept of lensing bands. Lensing bands have been explored in [98, 99] as powerful probes of new physics beyond GR, because they marginalize over astrophysical uncertainties.
@@
More specifically, a black-hole image always depends on (i) the spacetime geometry, (ii) the propagation equation for light and (iii) the distribution and properties of matter that emits and absorbs light. We are interested in testing (i) and (ii) independently of (iii), because our understanding of the astrophysical properties of accretion disks is not well enough developed to uniquely specify the emission and absorption profile of the accretion disk of a given black hole [88, 89, 90, 31, 100]. In addition, accretion disks are not always in a stationary state, so that time-dependence of the emission is in general expected [31, 101, 100, 102].
@@
There are two ways of addressing this uncertainty, both of which rely on photon rings. Photon rings arise because of strong gravitational lensing around a black hole. They constitute lensed images of the black hole which are present in addition to the direct image and which are exponentially stacked in GR. The $n$th photon ring in a black-hole image is constituted by those geodesics that, roughly speaking, complete $n$ half-orbits about the black hole, see [103] for a discussion of how to define a half-orbit in and beyond GR and [104, 105, 106, 107, 108, 109] for a discussion of photon rings. It can be shown that the direct emission that arrives on an observer’s screen from the vicinity of a black hole (the $n=0$ emission) depends on where in the spacetime the light was emitted, i.e., the location of the $n=0$ emission in the image is strongly dependent on the astrophysical properties of the accretion disk. For $n=1$, the dependence is already lower, as light that completes one half-orbit about the black hole before arriving at the observer’s screen necessarily arrives within a bounded region in the image, known as the $n=1$ photon ring. As $n$ increases, the size of this bounded region decreases and is therefore less and less dependent on the properties of the accretion disk. For $n\rightarrow\infty$, no dependence on the properties of the accretion disk is left. Therefore, photon rings constitute observables that probe spacetime geometry largely independently of astrophysical properties of the accretion disk, see, e.g., [108, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120].
@@
However, the $n\rightarrow\infty$ photon ring is not observable, because it has vanishing intensity. In practice, the $n=1$ photon ring may be detectable by future VLBI arrays, see, e.g., [71, 121, 116, 122]. Imposing an informed prior on the image reconstruction for M87*, it has been shown in [123] that a persistent (across consecutive observational days) thin ring is statistically favored in addition to a broad emission region over reconstructed images without a thin ring. Thin photon rings are generic in GR and beyond, thus motivating this choice of prior. This thin ring may be interpreted as the $n=1$ photon ring, but its reconstruction is of course predicated on the choice of prior.
@@
The location of the $n=1$ photon ring in the image depends on all three aspects (i-iii), i.e. the spacetime and light propagation properties as well as the properties of the accretion disk. Therefore, we calculate the lensing band, which is the region in the image within which the location of the $n=1$ photon ring varies when the emission region is varied in-between the black-hole horizon and infinitely far away. By definition, the $n=1$ photon ring always lies within the $n=1$ lensing band.
@@
As highlighted in [99], if the $n=1$ lensing band in a theory beyond GR does not overlap with the corresponding lensing band in GR for a given range of couplings, an $n=1$ photon ring detection would either rule out this range of couplings or rule out GR, with a confidence level that depends on the uncertainties in the photon ring detection.
@@ IV.2 Numerical method
Figure 1: We show the $n=1$ lensing band (which, for near edge-on inclinations, fills nearly the entire lower-half of the image) for the PPL polarization for three different inclinations $i$ and distinct coupling strengths $\alpha/M^{2}$. The lensing band for the Schwarzschild metric without non-minimal coupling is shown in black, and the lensing band with non-minimal coupling is overlaid in light pink. Wherever these regions overlap, the overlap is indicated in reddish brown.
We calculate lensing bands with the Flexible Object Oriented Ray Tracer FOORT [124], which ray-traces null geodesics in arbitrary spacetime geometries. FOORT has been previously used to explore image features of compact objects in modified theories of gravity [125, 113, 65].
@@
Figure 2: Area of the overlap between the $n=1$ lensing bands with non-minimal coupling and GR in terms of the coupling strength $\alpha/M^{2}$, for four different inclinations $i$. The total field of view of the image is $20M\times 20M$.
@@ V Results
Figure 3: Ratio of the area of the $n=1$ lensing bands with non-minimal light-curvature couplings (LCC) and GR in terms of the coupling strength $\alpha/M^{2}$ for four different inclinations $i$. The total area of the image is $20M\times 20M$.
@@
For inclinations $i\gg 0\degree$, the overlap between lensing bands with and without non-minimal coupling depends on the angle in the image: for instance for $\alpha/M^{2}=0.4$ and $i=50\degree$, there is no overlap between the lensing bands in the upper half of the image, whereas there is an overlapping area in the lower part of the image. This is because, as the inclination increases, the upper part of the lensing band becomes thinner, whereas the lower part stretches out until it covers the entire lower half of the image.
To diagnose in a more quantitative way how large the overlap between lensing bands is and to find the critical values of the coupling (both positive and negative), when no overlap exists, we calculate the overlap area in the following, for both polarizations.
@@ V.1 Quantitative comparison of lensing bands to the case without non-minimal coupling
In Fig. 2, we plot the overlap area between the $n=1$ lensing bands of the Schwarzschild black hole with and without the non-minimal coupling for both polarizations, normalized to the area of the $n=1$ lensing band of a Schwarzschild black hole without the non-minimal coupling (denoted as $\rm GR$ in the images). The two polarizations (PPL and PPM) strongly differ only at large inclination, i.e., close to edge-on. For both polarizations, the area of overlap with the $n=1$ Schwarzschild lensing band increases for $\alpha/M^{2}<0$, reaches a maximum at $\alpha/M^{2}=0$ (where the non-minimal coupling vanishes and we are back to GR) and decreases for $\alpha/M^{2}>0$ (except for $i=89^{\circ}$ where PPL has a plateau for $\alpha/M^{2}<0$ and PPM a plateau for $\alpha/M^{2}>0$). For both polarizations, there are sufficiently large values of $\alpha/M^{2}$ for which there is no overlap: $\alpha/M^{2}\gtrsim 0.6$ at $i=1^{\circ}$ and $\alpha/M^{2}\gtrsim 0.9$ at $i=17^{\circ}$ (inclination of M87*). Those large values of $\alpha$ delineate a regime where the lensed photons for GR and for a non-minimal coupling are contained within fully distinct bands on the observer’s screen. To exclude the corresponding values of couplings based on observations, it is necessary to combine EHT-observations with an independent mass measurement that sets the expected diameter of the photon ring in the image plane. Without such an independent mass measurement, a rescaling of the mass can always bring the photon ring diameter with non-minimal coupling into agreement with a photon ring without non-minimal coupling and at appropriately rescaled mass.
@@
In Fig. 3 we compare the absolute areas of the $n=1$ lensing bands on the image plane with those of GR. At each inclination there is a crossing between PPL and PPM at the GR value $\alpha/M^{2}=0$: PPL dominates for $\alpha/M^{2}<0$ and PPM takes over for $\alpha/M^{2}>0$. Hence, the area spanned by the lensing bands with non-minimal couplings is larger (smaller) than in GR as long as $\alpha/M^{2}<0$ and smaller (larger) than in GR for $\alpha/M^{2}>0$ for PPL (respectively, PPM). We conclude that low inclinations are more favorable to constrain negative values of $\alpha$, because there is a more pronounced discrepancy between the two polarizations.
@@
In Fig. 4 we compare the area of the $n=1$ lensing bands of the two distinct polarizations. We observe that $A_{\rm PPM}<A_{\rm PPL}$ for $\alpha/M^{2}<0$, while $A_{\rm PPM}>A_{\rm PPL}$ for $\alpha/M^{2}>0$, with very weak dependence in the inclination, in agreement with Fig. 3. More interestingly, for low-enough inclinations $i\lesssim 17^{\circ}$, there are intervals $-1/2<\alpha/M^{2}\leq-0.4$ and $10.4\leq\alpha/M^{2}<$ 1 for which there is no overlap between the lensing bands of the two polarizations. This is in strong contrast to the case without non-minimal coupling, where the $n=1$ lensing band has no polarization dependence. It also opens an intriguing observational possibility, because the EHT actually reports information on polarization, see, e.g., [23, 26, 33, 34]. Therefore, a reconstruction of the $n=1$ photon ring in two polarizations corresponding to PPL and PPM constrains $\alpha$ without the need for an independent measurement of the mass of the black hole.
As one would expect, the area of overlap between the lensing bands of the two polarizations quickly grows as the GR case $\alpha/M^{2}=0$ is approached and only coupling values $\alpha\sim\mathcal{O}(M^{2})$ can be constrained. Also, at high inclination $i=89^{\circ}$, the overlap areas are similar for almost all values of $\alpha/M^{2}$ as $A_{\rm PPM}\sim A_{\rm PPL}$, hence the two polarizations are almost indistinguishable. In this context, it is interesting that the inclination for M87* is $i_{\mathrm{M}87^{*}}\approx 17\,\degree$, which can in principle provide significantly stronger constraints.
@@
Figure 4: Top: Ratio of the area of the $n=1$ lensing bands for the PPM and PPL polarizations in terms of the coupling strength $\alpha/M^{2}$, for four different inclinations $i$. Bottom: Ratio of the area of the overlap between the $n=1$ lensing bands for the PPM and PPL polarizations and the area of the PPL $n=1$ lensing band in terms of the coupling strength $\alpha/M^{2}$, for four different inclinations $i$. The total area of the image is $20M\times 20M$.
@@ V.2 A case study: the M87* photon ring
By imposing an informed prior on the image reconstruction of M87*, Ref. [123] demonstrated that a persistent thin ring, visible across multiple observational days, is statistically preferred alongside a broader emission region when compared to reconstructions lacking such a ring. Thin photon rings generically appear in shadow images of both GR and beyond-GR (ultra-)compact objects [126, 127, 61, 62, 128, 129, 77], motivating this choice of prior. While the reconstructed thin ring may be interpreted as the $n=1$ photon ring, its presence ultimately depends on the adopted prior. As a case study to explore whether future VLBI-arrays can, in principle, constrain the non-minimal couplings we consider in this work, we take the results of Ref. [123] at face value. That is, we assume that the thin ring matches the $n=1$ photon ring of M87* which, due to the low inclination $i_{\rm M87^{*}}\approx 17^{\circ}$ of M87*, is well-approximated by a circle.
In this case, the angular radius of the M87* photon ring is $\theta_{\rm ring}=21.74\pm 0.10\,\mu{\rm as}$ [123], the distance to M87* is $d=16.8\pm 0.8\,{\rm Mpc}$ [22], while external mass measurements from gas dynamics result in $M_{\rm M87^{*}}=(6.60\pm 0.40)\cdot 10^{9}{\rm M_{\odot}}$ [130]. Using these values, we can obtain the diameter $x_{\rm PR}/M_{\rm M87^{*}}$ of the photon ring in image-plane coordinates via
@@
At the inclination $i=17^{\circ}$ and for the whole range of couplings, we take a cross section of our images at $y=0$ and compare the region along the $x$-axis where the $n=1$ photon ring with non-minimal coupling can be located with the value for M87* given in Eq. (16). The results are shown in Fig. 5, from which we see that the only values of $\alpha/M^{2}_{\rm M87^{*}}$ compatible with observations are $-0.3\lesssim\alpha/M^{2}_{\rm M87^{*}}\lesssim 0.3$. These constraints result in the upper bound
@@
This is also the range where PPL and PPM lensing bands overlap along the $x$-axis.
Figure 5: In the image plane, we fix $y=0$ and compare the allowed range of $x_{\rm PR}$ where the $n=1$ photon ring can be located (i.e., the lensing band) for the cases of Schwarzschild (gray), PPL (red) and PPM (blue) with the measured value (in green), assuming an inclination of $i=17^{\circ}=i_{\rm M87^{*}}$.

### v2

@@ IV.1 Background on lensing bands
To explore how the non-minimal coupling $\alpha$ affects VLBI observations of supermassive black holes, we leverage the concept of lensing bands. Lensing bands have been explored in [Paugnat:2022qzy, Cardenas-Avendano:2023obg] as powerful probes of new physics beyond GR, because they marginalize over astrophysical uncertainties.
@@
More specifically, a black-hole image always depends on (i) the spacetime geometry, (ii) the propagation equation for light and (iii) the distribution and properties of matter that emits and absorbs light. We are interested in testing (i) and (ii) independently of (iii), because our understanding of the astrophysical properties of accretion disks is not well enough developed to uniquely specify the emission and absorption profile of the accretion disk of a given black hole [Lara:2021zth, Ozel:2021ayr, Kocherlakota:2022jnz, EventHorizonTelescope:2022urf, Chatterjee:2022pbo]. In addition, accretion disks are not always in a stationary state, so that time-dependence of the emission is in general expected [EventHorizonTelescope:2022urf, Murchikova:2022aiz, Chatterjee:2022pbo, Blaes:2025usm].
@@
There are two ways of addressing this uncertainty, both of which rely on photon rings. Photon rings arise because of strong gravitational lensing around a black hole. They constitute lensed images of the black hole which are present in addition to the direct image and which are exponentially stacked in GR. The $n$th photon ring in a black-hole image is constituted by those geodesics that, roughly speaking, complete $n$ half-orbits about the black hole, see [Eichhorn:2022bbn] for a discussion of how to define a half-orbit in and beyond GR and [Johnson:2019ljv, Gralla:2019xty, Gralla:2020srx, Gralla:2020yvo, Broderick:2021ohx, Chael:2021rjo] for a discussion of photon rings. It can be shown that the direct emission that arrives on an observer’s screen from the vicinity of a black hole (the $n=0$ emission) depends on where in the spacetime the light was emitted, i.e., the location of the $n=0$ emission in the image is strongly dependent on the astrophysical properties of the accretion disk. For $n=1$, the dependence is already lower, as light that completes one half-orbit about the black hole before arriving at the observer’s screen necessarily arrives within a bounded region in the image, known as the $n=1$ photon ring. As $n$ increases, the size of this bounded region decreases and is therefore less and less dependent on the properties of the accretion disk. For $n\rightarrow\infty$, no dependence on the properties of the accretion disk is left. Therefore, photon rings constitute observables that probe spacetime geometry largely independently of astrophysical properties of the accretion disk, see, e.g., [Broderick:2021ohx, Wielgus:2021peu, Eichhorn:2022oma, Ayzenberg:2022twz, Staelens:2023jgr, Kocherlakota:2023qgo, Aratore:2024bro, Galison:2024bop, KumarWalia:2024yxn, Farah:2024mkq, KumarWalia:2024omf, Keeble:2025gbj].
@@
However, the $n\rightarrow\infty$ photon ring is not observable, because it has vanishing intensity. In practice, the $n=1$ photon ring may be detectable by future VLBI-arrays, see, e.g., [Ayzenberg:2023hfw, Lupsasca:2024xhq, Galison:2024bop, Tamar:2024tfc]. Imposing an informed prior on the image reconstruction for M87*, it has been shown in [Broderick:2022tfu] that a persistent (across consecutive observational days) thin ring is statistically favored in addition to a broad emission region over reconstructed images without a thin ring. Thin photon rings are generic in GR and beyond, thus motivating this choice of prior. This thin ring may be interpreted as the $n=1$ photon ring, but its reconstruction is of course predicated on the choice of prior.
@@
The location of the $n=1$ photon ring in the image depends on all three aspects (i-iii), i.e. the spacetime and light propagation properties as well as the properties of the accretion disk. However, the $n=1$ photon ring must be contained within a bounded region in the image, known as the $n=1$ lensing band. The $n=1$ lensing band is defined as the set of points on the observer’s screen to which at least one photon completing one half-orbit about the black hole arrives after being emitted at an arbitrary location on the equatorial plane. Accordingly, the lensing band can be understood as the region in the image within which the location of the $n=1$ photon ring varies when the emission region is varied in-between the black-hole horizon and infinitely far away. By definition, the $n=1$ photon ring always lies within the $n=1$ lensing band.
@@
As highlighted in [Cardenas-Avendano:2023obg], if the $n=1$ lensing band in a theory beyond GR does not overlap with the corresponding lensing band in GR for a given range of couplings, an $n=1$ photon ring detection would either rule out this range of couplings or rule out GR, with a confidence level that depends on the uncertainties in the photon ring detection.
@@ IV.2 Numerical method
Figure 1: We show the $n=1$ lensing band (which, for near edge-on inclinations, fills nearly the entire lower-half of the image) for the PPL polarization for three different inclinations $i$ and distinct coupling strengths $\alpha/M^{2}$. The lensing band for the Schwarzschild metric without non-minimal coupling is shown in black, and the lensing band with non-minimal coupling is overlaid in light pink. Wherever these regions overlap, the overlap is indicated in reddish brown. The lensing bands expand inwards (outwards) for negative (positive) value of the coupling. This behaviour is reversed for the PPM polarization.
We calculate lensing bands with the Flexible Object Oriented Ray Tracer FOORT [Mayerson:2025xxx], which ray-traces null geodesics in arbitrary spacetime geometries. FOORT has been previously used to explore image features of compact objects in modified theories of gravity [Mayerson:2023wck, Staelens:2023jgr, Fernandes:2024ztk].
@@
Figure 2: Area of the overlap between the $n=1$ lensing bands with non-minimal coupling and GR (reddish brown regions in Fig. 1 for the PPL case) in terms of the coupling strength $\alpha/M^{2}$, for four different inclinations ii. The total field of view of the image is $20M\times 20M$. For all inclinations, generically, the PPL lensing bands have a larger overlap with the GR lensing band, compared to the PPM case, for negative coupling, and vice-versa for positive couplings. For large inclinations, the overlap area can plateau because the lower-half of the image is filled by the lensing band both in GR and with non-minimal couplings.
@@ V Results
Figure 3: Ratio of the total area of the $n=1$ lensing bands with non-minimal light-curvature couplings (LCC) and GR in terms of the coupling strength $\alpha/M^{2}$ for four different inclinations ii. This figure compares how much larger or smaller, compared to GR, is the region in the image plane where the $n=1$ photon ring can be located. The total area of the image is $20M\times 20M$. For negative values of the coupling, there is a larger disparity between the two polarizations for small inclinations, making them more favorable to constrain the coupling.
@@
For inclinations $i\gg 0\degree$, the overlap between lensing bands with and without non-minimal coupling depends on the angle in the image: for instance, for $\alpha/M^{2}=0.4$ and $i=50\degree$, there is no overlap between lensing bands in the upper-half of the image, whereas there is an overlapping area in the lower part of the image. This is because, as the inclination increases, the upper part of the lensing band becomes thinner, whereas the lower part stretches out until it covers the entire lower-half of the image.
To diagnose in a more quantitative way how large the overlap between lensing bands is and to find the critical values of the coupling (both positive and negative) when no overlap exists, we calculate the overlap area in the following, for both polarizations.
@@ V.1 Quantitative comparison of lensing bands to the case without non-minimal coupling
In Fig. 2 we plot the overlap area between the $n=1$ lensing bands of the Schwarzschild black hole with and without the non-minimal coupling for both polarizations, normalized to the area of the $n=1$ lensing band of a Schwarzschild black hole without the non-minimal coupling (denoted as $\rm GR$ in the images). The two polarizations (PPL and PPM) strongly differ only at large inclination, i.e., close to edge-on. For both polarizations, the area of overlap with the $n=1$ Schwarzschild lensing band increases for $\alpha/M^{2}<0$, reaches a maximum at $\alpha/M^{2}=0$ (where the non-minimal coupling vanishes and we are back to GR) and decreases for $\alpha/M^{2}>0$ (except for $i=89^{\circ}$ where PPL has a plateau for $\alpha/M^{2}<0$ and PPM a plateau for $\alpha/M^{2}>0$). For both polarizations, there are sufficiently large values of $\alpha/M^{2}$ for which there is no overlap: $\alpha/M^{2}\gtrsim 0.6$ at $i=1^{\circ}$ and $\alpha/M^{2}\gtrsim 0.9$ at $i=17^{\circ}$ (inclination of M87*). Those large values of $\alpha$ delineate a regime where the lensed photons for GR and for a non-minimal coupling are contained within fully distinct bands on the observer’s screen. To exclude the corresponding values of couplings based on observations, it is necessary to combine EHT-observations with an independent mass measurement that sets the expected diameter of the photon ring in the image plane. Without such an independent mass measurement, a rescaling of the mass can always bring the photon ring diameter with non-minimal coupling into agreement with a photon ring without non-minimal coupling and at appropriately rescaled mass.
@@
In Fig. 3 we compare the absolute areas on the image plane of the $n=1$ lensing bands with non-minimal coupling with those of GR. At each inclination, there is a crossing between PPL and PPM at the GR value $\alpha/M^{2}=0$: PPL dominates for $\alpha/M^{2}<0$ and PPM takes over for $\alpha/M^{2}>0$. Hence, the area spanned by the lensing bands with non-minimal couplings is larger (smaller) than in GR as long as $\alpha/M^{2}<0$ and smaller (larger) than in GR for $\alpha/M^{2}>0$ for PPL (respectively, PPM).
@@
In Fig. 4 we compare the area of the $n=1$ lensing bands of the two distinct polarizations. We observe that $A_{\rm PPM}<A_{\rm PPL}$ for $\alpha/M^{2}<0$, while $A_{\rm PPM}>A_{\rm PPL}$ for $\alpha/M^{2}>0$, with very weak dependence in the inclination, in agreement with Fig. 3. More interestingly, for low-enough inclinations $i\lesssim 17^{\circ}$, there are intervals $1/2<\alpha/M^{2}\leq-0.4$ and $0.4\leq\alpha/M^{2}<1$ for which there is no overlap between the lensing bands of the two polarizations. This is in strong contrast to the case without non-minimal coupling, where the $n=1$ lensing band has no polarization dependence. It also opens an intriguing observational possibility, because the EHT actually reports information on polarization, see, e.g., [EventHorizonTelescope:2021bee, EventHorizonTelescope:2023gtd, EventHorizonTelescope:2024hpu, EventHorizonTelescope:2024rju]. Therefore, a reconstruction of the $n=1$ photon ring in two polarizations corresponding to PPL and PPM constrains $\alpha$ without the need for an independent measurement of the mass of the black hole.
As one would expect, the area of overlap between the lensing bands of the two polarizations quickly grows as the GR case $\alpha/M^{2}=0$ is approached and only coupling values $\alpha\sim\mathcal{O}(M^{2})$ can be constrained. Also, at high inclination $i=89^{\circ}$, the overlap areas are similar for almost all values of $\alpha/M^{2}$ as $A_{\rm PPM}\sim A_{\rm PPL}$ due to the enlargement of the PPL and PPM lensing bands in the lower-half plane caused by relativistic effects. On the other hand, for high inclination angles there is no overlap in the upper-half plane of the image, as seen in Fig. 5, and overall the lensing bands associated with both polarizations may remain distinguishable. These differences can be captured quantitatively by calculating the ratio between the lensing bands of the two polarizations restricted to the upper-half plane of the image, which we plot in Fig. 6. We caution that actual observational data, with sparse coverage in the Fourier plane, may be much more difficult to interpret in terms of such observables. Nevertheless, our results indicate that it may be relevant to develop observables that are adapted to only parts of the resulting image, such as, in our case, only the upper half plane. We believe this is worth exploring in detail elsewhere.
@@
Figure 4: Top: Ratio of the area of the $n=1$ lensing bands for the PPM and PPL polarizations in terms of the coupling strength $\alpha/M^{2}$, for four different inclinations ii. Bottom: Ratio of the area of the overlap between the $n=1$ lensing bands for the PPM and PPL polarizations and the area of the PPL $n=1$ lensing band in terms of the coupling strength $\alpha/M^{2}$, for four different inclinations ii. The total area of the image is $20M\times 20M$. For large inclinations, the overlap area plateaus because both polarizations fill a large portion of the lower-half plane of the image, and no overlap exists in the upper-half plane of the image. For sufficiently small inclinations, and large absolute values of the coupling, there is no overlap between the lensing bands, as increasing $|\alpha|$ pushes them in different directions.
Figure 5: We show the $n=1$ lensing bands (which, for near edge-on inclinations, nearly fill the entire lower-half plane of the image) for the PPL and PPM polarizations for the largest inclination $i=89^{\circ}$ and coupling strength $\alpha/M^{2}=0.9$. The total area of the image is $20M\times 20M$. The lensing band for the PPL polarization is shown in light pink, and the lensing band for the PPM polarization is overlaid in blue. The region where these lensing bands overlap is indicated in purple. In the upper-half plane $y\geq 0$, the PPL lensing band is everywhere located outward from the PPM lensing band, while the PPL and PPM lensing bands greatly overlap in the lower-half plane $y<0$, filling it almost entirely.
Figure 6: Top: Ratio of the area of the $n=1$ lensing bands for the PPM and PPL polarizations in terms of the coupling strength $\alpha/M^{2}$, for four different inclinations $i$, considering only the upper-half plane $y\geq 0$. Bottom: Ratio of the area of the overlap between the $n=1$ lensing bands for the PPM and PPL polarizations and the area of the PPL $n=1$ lensing band in terms of the coupling strength $\alpha/M^{2}$, for four different inclinations $i$, considering only the upper-half plane $y\geq 0$. The total area of the image is $10M\times 20M$. The overlap area decreases with increasing inclination as the PPL and PPM lensing bands are clearly distinct in the upper-half plane at $i=89^{\circ}$, see Fig. 5.
@@ V.2 A case study: the M87* photon ring
By imposing an informed prior on the image reconstruction of M87*, Ref. [Broderick:2022tfu] demonstrated that a persistent thin ring, visible across multiple observational days, is statistically preferred alongside a broader emission region when compared to reconstructions lacking such a ring. Thin photon rings generically appear in shadow images of both GR and beyond-GR (ultra-)compact objects [Guerrero:2021ues, Olmo:2021piq, Eichhorn:2022fcl, Carballo-Rubio:2022aed, Guerrero:2022msp, daSilva:2023jxa, Carballo-Rubio:2025fnc], motivating this choice of prior. While the reconstructed thin ring may be interpreted as the $n=1$ photon ring, its presence ultimately depends on the adopted prior. As a case study to explore whether future VLBI-arrays can, in principle, constrain the non-minimal coupling we consider in this work, we take the results of Ref. [Broderick:2022tfu] at face value. That is, we assume that the thin ring matches the $n=1$ photon ring of M87* which, due to the low inclination $i_{\rm M87^{*}}\approx 17^{\circ}$ of M87*, is well-approximated by a circle.
In this case, the angular radius of the M87* photon ring is $\theta_{\rm ring}=21.74\pm 0.10\,\mu{\rm as}$ [Broderick:2022tfu], the distance to M87* is $d=16.8\pm 0.8\,{\rm Mpc}$ [EventHorizonTelescope:2019ggy], while external mass measurements from gas dynamics result in $M_{\rm M87^{*}}=(6.60\pm 0.40)\cdot 10^{9}{\rm M_{\odot}}$ [Gebhardt:2011yw]. Using these values, we can obtain the dimensionless diameter $x_{\rm PR}/M_{\rm M87^{*}}$ of the photon ring in image-plane coordinates via
@@
At the inclination $i=17^{\circ}$ and for the whole range of couplings, we take a cross section of our images at $y=0$ and compare the region along the $x$-axis where the $n=1$ photon ring with non-minimal coupling can be located with the value for M87* given in Eq. (16). The results are shown in Fig. 7, from which we see that the only values of $\alpha/M^{2}_{\rm M87^{*}}$ compatible with unpolarized observations – for which PPL and PPM lensing bands overlap along the $x$-axis – are $0.3\lesssim\alpha/M^{2}_{\rm M87^{*}}\lesssim 0.3$. [^7] Most of the analyses of the bright thin ring apparent in shadow images of M87* have been performed by the EHT Collaboration with unpolarized light, as the linearly- and circularly-polarized components of the lensed radiation captured by the current VLBI-array are found to be subdominant, but have nevertheless been successfully extracted [EventHorizonTelescope:2021bee, EventHorizonTelescope:2021srq, EventHorizonTelescope:2023gtd, EventHorizonTelescope:2025vum]. These constraints result in the upper bound
@@
Figure 7: In the image plane, we fix $y=0$ (see the lensing bands in Fig. 1) and compare the allowed range of $x_{\rm PR}$ where the $n=1$ photon ring can be located (i.e., the lensing band) for the cases of Schwarzschild (gray), PPL (red) and PPM (blue) with the measured value (in green), assuming an inclination of $i=17^{\circ}=i_{\rm M87^{*}}$. Regions where there is no overlap between the PPL and PPM cases are excluded as they are not compatible with unpolarized observations.